### v1

# Online Bayesian Experimental Design for Partially Observable Dynamical Systems

Abstract
Bayesian experimental design (BED) provides a principled framework for optimizing data collection, but existing approaches do not apply to crucial real-world settings such as dynamical systems with partial observability, where only noisy and incomplete observations are available. These systems are naturally modeled as state-space models (SSMs), where latent states mediate the link between parameters and data, making the likelihood—and thus information-theoretic objectives like the expected information gain (EIG)—intractable. In addition, the dynamical nature of the system requires online algorithms that update posterior distributions and select designs sequentially in a computationally efficient manner. We address these challenges by deriving new estimators of the EIG and its gradient that explicitly marginalize latent states, enabling scalable stochastic optimization in nonlinear SSMs. Our approach leverages nested particle filters (NPFs) for efficient online inference with convergence guarantees. Applications to realistic models, such as the susceptible–infected–recovered (SIR) and a moving source location task, show that our framework successfully handles both partial observability and online computation.
1 INTRODUCTION
Designing experiments that yield maximally informative data is a fundamental problem across science and engineering. The ability to guide data collection efficiently can accelerate learning in diverse domains, including material discovery (lei2021bayesian; lookman2019active), DNA sequencing (weilguny2023dynamic), sensor networks (wu2023offline), and drug discovery (lyu2019ultra). Formally, given a known model structure, let $\boldsymbol{\theta}$ represent the parameters of interest and $\boldsymbol{\xi}$ the design chosen for the experiment, which determines the corresponding observation $\boldsymbol{y}$. The objective is to select designs that maximize the information gained about $\boldsymbol{\theta}$, thereby reducing uncertainty as more data becomes available.
Bayesian experimental design (BED) provides a principled framework for combining Bayesian inference with data acquisition (lindley1956measure; huan2024optimal). In this setting, prior beliefs about the parameters $p(\boldsymbol{\theta})$ are updated as data are collected, yielding a posterior distribution that quantifies uncertainty. This update is driven by the likelihood $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, which links parameters to observations under a given design $\boldsymbol{\xi}$. When this process is repeated adaptively, using past observations to guide future design choices, we obtain the sequential setting known as Bayesian adaptive design (BAD) (Rainforth24). At the experiment $t$, the prior is $p(\boldsymbol{\theta}\!\mid\!h_{t-1})$, where $h_{t-1}=\{\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1}\}$ is the history of past designs and observations. The next design $\boldsymbol{\xi}_{t}$ is then chosen to maximize some utility function, commonly the expected information gain (EIG), which requires evaluating $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$.
A common assumption in BED is that the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ is tractable and available in closed form. While this holds in some settings, it is unrealistic in many real-world scenarios where the system cannot be fully observed. In particular, dynamical systems often involve latent states that evolve over time and are only observed through noisy, partial measurements. Such systems are naturally modeled as state-space models (SSMs) (Sarkka23), where the likelihood is defined only indirectly through integration over the hidden states, and is therefore intractable.
Several BED approaches address implicit likelihoods, i.e., cases where we can simulate observations from the model but cannot evaluate $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ in closed form. In principle, such methods could also be applied or extended to SSMs, since they provide a way to simulate samples. Most existing BED approaches rely on amortization, where a surrogate (e.g., a neural network) is trained offline on a fixed dataset to approximate the likelihood, and then reused at deployment. Such methods have been proposed, mostly in static BED settings  (Kleinegesse20; Dehideniya18), with recent extensions to sequential settings via mutual-information bounds (Ivanova21) or density-ratio estimation (Kleinegesse21).
In contrast, we focus on the non-amortized setting: designs are optimized sequentially, and their accuracy can be improved arbitrarily by allocating more computational resources at each step. This setting is particularly relevant in applications where experiments themselves are costly, such as adaptive clinical trials (chaloner1995bayesian) and astronomical observations (loredo2004bayesian). In such cases, the cost of additional computation between measurements is negligible compared to the cost of data collection. Our aim is therefore to develop methods that retain full sequential adaptability while ensuring that design optimization at each step remains as accurate as resources allow. We provide further discussion of related work in Section 4.
Contributions. In this paper, we propose a novel framework for BAD in partially observable dynamical systems, where likelihoods are intractable and online inference is required. Our main contributions are (i) new estimators of EIG and its gradient, specifically derived for state-space models with latent states, enabling design optimization under partial observability, and (ii) an online BAD approach based on nested particle filters (NPFs) (Crisan18bernoulli), which recursively approximates the joint posterior over parameters and latent states without reprocessing the full history. In representative experiments, we show that our framework successfully enables BED under partial observability and online constraints.
2 PROBLEM SETTING
2.1 Bayesian adaptive design (BAD)
The goal of BAD is to select, at each experiment $t$, a design $\boldsymbol{\xi}_{t}$ that maximizes the EIG about the parameters:
$\boldsymbol{\xi}_{t}^{\star}=\arg\max_{\boldsymbol{\xi}_{t}\in\Omega}\;\mathcal{I}(\boldsymbol{\xi}_{t}),$
where $\Omega$ is the design space.
The EIG quantifies the reduction in posterior uncertainty about $\boldsymbol{\theta}$ after observing $\boldsymbol{y}_{t}$ under design $\boldsymbol{\xi}_{t}$:
$\mathcal{I}$ $(\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{y}_{t}\mid\boldsymbol{\xi}_{t})}\!\bigg[\,\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t-1})]-\mathcal{H}[p(\boldsymbol{\theta}\!\mid\!h_{t})]\bigg]$
$=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})\,p(\boldsymbol{y}_{t}\mid\boldsymbol{\theta},\boldsymbol{\xi}_{t})}\!\big[\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})-\log p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})\big],$ (1)
with $h_{t-1}=(\boldsymbol{\xi}_{1:t-1},\boldsymbol{y}_{1:t-1})$ the history, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ the likelihood, $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})$ the marginal predictive distribution, and $\mathcal{H}$ denoting entropy.
In many applications, however, observations $\boldsymbol{y}_{t}$ are not generated directly by $\boldsymbol{\theta}$, but through an underlying dynamical system with latent states $\boldsymbol{x}_{0:t}$. Such cases are naturally modeled as SSMs, resulting in a setting where the likelihood is not available in closed form, rendering standard BAD methods inapplicable.
2.2 State-space models (SSMs)
We focus on dynamical systems that can be described by Markovian SSMs. At each experiment or time step $t\in\mathbb{N}$, the system state $\boldsymbol{x}_{t}\in\mathbb{R}^{d_{x}}$ evolves and produces an observation $\boldsymbol{y}_{t}\in\mathbb{R}^{d_{y}}$:
$\boldsymbol{x}_{t}$ $\sim$ $f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1},\boldsymbol{\theta},\boldsymbol{\xi}_{t}),$ (2)
$\boldsymbol{y}_{t}$ $\sim$ $g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t}),$ (3)
where $\boldsymbol{\theta}\in\mathbb{R}^{d_{\theta}}$ are unknown parameters and $\boldsymbol{\xi}_{t}\in\Omega\subset\mathbb{R}^{d_{\xi}}$ is the design variable. Here, $f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1},\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ is the transition density of the latent states, and $g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ is the observation model, both assumed to be differentiable with respect to (w.r.t.) $\boldsymbol{\xi}_{t}$. We assume that $\boldsymbol{y}_{t}$ (conditional on the state, parameters, and design) is conditionally independent of all other observations, and the prior probability density functions (pdfs) of the state and the parameters, $p(\boldsymbol{\theta})$ and $p(\boldsymbol{x}_{0})$, are known and independent.
2.3 Particle filters (PFs)
In SSMs, the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t},h_{t-1})$ is defined only through marginalization over latent states, which is intractable in general. Particle filters (PFs), or sequential Monte Carlo methods (Gordon93; Doucet00; Doucet01; Djuric03), provide a standard solution by recursively approximating the filtering or posterior distribution, by representing it as a weighted set of $M$ particles,
$p(\mathrm{d}\boldsymbol{x}_{t}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\theta},\boldsymbol{\xi}_{1:t})\approx\sum_{m=1}^{M}w_{\boldsymbol{x},t}^{(m)}\,\delta_{\boldsymbol{x}_{t}^{(m)}}(\mathrm{d}\boldsymbol{x}_{t}).$
At each step, particles $\boldsymbol{x}_{t}^{(m)}$ are propagated forward using the state transition $f$, while their weights $w_{\boldsymbol{x},t}^{(m)}$ are updated in proportion to the observation model $g$. This allows the algorithm to adaptively focus on regions of high likelihood and provides a flexible representation of complex, nonlinear distributions.
Standard PFs assume that both the parameters $\boldsymbol{\theta}$ and the designs $\boldsymbol{\xi}_{1:t}$ are known and fixed, so that only the latent states need to be inferred. In BAD, however, the parameters $\boldsymbol{\theta}$ are also unknown and must be estimated. This motivates the use of NPFs, which extend the PF framework to incorporate parameter inference.
2.4 Nested particle filters (NPFs)
NPFs (Crisan18bernoulli) extend PFs to jointly estimate states and parameters in an online manner, targeting the joint posterior $p(\mathrm{d}\boldsymbol{x}_{0:t},\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. An NPF employs two intertwined layers of PFs: (i) an inner layer that estimates the state distribution conditional on each parameter particle
$p(\mathrm{d}\boldsymbol{x}_{t}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\theta},\boldsymbol{\xi}_{1:t})\approx\sum_{m=1}^{M}\sum_{n=1}^{N}w_{\boldsymbol{\theta},t}^{(m)}w_{\boldsymbol{x},t}^{(m,n)}\,\delta_{\boldsymbol{x}_{t}^{(m,n)}}(\mathrm{d}\boldsymbol{x}_{t}),$
and (ii) an outer layer that represents the parameter posterior
$p(\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})\approx\sum_{m=1}^{M}w_{\boldsymbol{\theta},t}^{(m)}\,\delta_{\boldsymbol{\theta}_{t}^{(m)}}(\mathrm{d}\boldsymbol{\theta}).$
This nested structure is related to sequential Monte Carlo squared (SMC2) (chopin2013smc2), but with a key computational improvement: instead of re-running inner filters from scratch, NPFs perturb parameters using a jittering kernel,
$\boldsymbol{\theta}_{t}^{(m)}\sim\kappa_{N}(\mathrm{d}\boldsymbol{\theta}\!\mid\!\boldsymbol{\theta}_{t-1}^{(m)}),$
so that the inner filters are updated a single step forward. This avoids reprocessing past data and yields linear cost $\mathcal{O}(t)$, compared to the quadratic $\mathcal{O}(t^{2})$ cost of SMC2.
Under suitable conditions, jittering prevents particle degeneracy and ensures consistency, providing asymptotic convergence guarantees (Crisan18bernoulli). This has motivated several extensions and further analysis, including Gaussian approximations (Perez-Vieites21) and Rao–Blackwellized variants (Fang23).
3 PROPOSED FRAMEWORK
We propose a new BAD framework for partially observable dynamical systems, deriving recursive estimators of the EIG and its gradient. This enables online optimization in continuous design spaces, $\boldsymbol{\xi}_{t}\in\mathbb{R}^{d_{\xi}}$, using stochastic gradient ascent (SGA). We further establish consistency of the proposed EIG estimator, and conclude the section by presenting the overall algorithmic scheme required for practical implementation.
3.1 EIG in SSMs
The EIG expression in (2.1) shows the dependency on the likelihood $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})$ and the marginal likelihood (evidence) $p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})}[\,p(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi}_{t})]$. Even when the likelihood is available in closed-form, estimating $\mathcal{I}(\boldsymbol{\xi}_{t})$ is doubly intractable because the evidence appears inside the outer expectation that defines the utility, typically needing nested Monte Carlo (NMC) methods or other approximations (Rainforth18; Foster19).
In partially observable dynamical systems, the difficulty compounds because both likelihood and evidence require marginalization over latent states. For compactness we denote the likelihood and the evidence as
$L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ $\coloneqq\mathbb{E}_{p(\boldsymbol{x}_{0:t}\mid\boldsymbol{\theta},h_{t-1})}\!\big[\,g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})\big],$
$Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ $\coloneqq\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})\,p(\boldsymbol{x}_{0:t}\mid\boldsymbol{\theta},h_{t-1})}\!\big[\,g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})\big],$
respectively. Substituting these into (2.1) yields
$\mathcal{I}(\boldsymbol{\xi}_{t})=\mathbb{E}_{p(\boldsymbol{\theta}\mid h_{t-1})\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\boldsymbol{y}_{t},\boldsymbol{x}_{0:t}\mid\boldsymbol{\theta},\boldsymbol{\xi}_{t})}}\left[\,\log\frac{L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}\right].$ (4)
This expression makes explicit that evaluating the EIG requires not only the usual outer expectation over $\boldsymbol{y}_{t}$ (and $\boldsymbol{\theta}$), but also inner expectations over latent-state trajectories inside both $L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}$ and $Z_{\boldsymbol{\xi}_{t}}$. Although these quantities are analytically intractable, they admit natural Monte Carlo approximations: $L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ can be estimated by sampling state trajectories conditional on a fixed $\boldsymbol{\theta}$, while $Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ requires sampling jointly from the parameter prior and state dynamics. This compounded intractability motivates the gradient representation and efficient Monte Carlo estimators that we develop next.
3.2 Gradient of the EIG
Using Fisher’s identity (Douc14) on the EIG of (4), we obtain a gradient representation that separates (i) derivatives of the likelihood and evidence and (ii) the design score at time $t$. Let
$\Gamma_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t},\boldsymbol{x}_{0:t},\boldsymbol{\theta})$ $\coloneqq g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1},\boldsymbol{\theta},\boldsymbol{\xi}_{t})\,\times$
$\quad\times p(\boldsymbol{x}_{0:t-1}\!\mid\!\boldsymbol{\theta},h_{t-1})p(\boldsymbol{\theta}\!\mid\!h_{t-1})$
be the joint distribution of states, parameters, and observations given a design (generative model at time $t$). We assume the availability of an approximation to the joint posterior distribution of states and parameters at the previous time step, $p(\boldsymbol{x}_{0:t-1},\boldsymbol{\theta}\!\mid\!h_{t-1})$.
The gradient of the EIG can be written as
$\nabla_{\boldsymbol{\xi}_{t}}\mathcal{I}(\boldsymbol{\xi}_{t})$ $=\mathbb{E}_{\Gamma_{\boldsymbol{\xi}_{t}}}\!\Bigg[\frac{\nabla_{\boldsymbol{\xi}_{t}}L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}-\frac{\nabla_{\boldsymbol{\xi}_{t}}Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}$
$~+\log\!\left[\,\frac{L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}{Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})}\right]s_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{x}_{t-1:t},\boldsymbol{y}_{t})\Bigg]$ (5)
where $\Gamma_{\boldsymbol{\xi}_{t}}$} stands for $\Gamma_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t},\boldsymbol{x}_{0:t},\boldsymbol{\theta})$}, and the design score at time $t$ is
$s_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{x}_{t-1:t},\boldsymbol{y}_{t})$ $\coloneqq\nabla_{\boldsymbol{\xi}_{t}}\log g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t})$
$\quad+\nabla_{\boldsymbol{\xi}_{t}}\log f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1},\boldsymbol{\theta},\boldsymbol{\xi}_{t}),$
with any of the terms vanishing when they do not depend on $\boldsymbol{\xi}_{t}$. See details in Appendix LABEL:supp-ap:gradient.
Direct evaluation of (3.2) is intractable. Both the likelihood $L_{\boldsymbol{\theta},\boldsymbol{x}_{t}}(\boldsymbol{y}_{t})$ and the evidence $Z_{\boldsymbol{x}_{t}}(\boldsymbol{y}_{t})$, as well as their gradients, are defined by expectations over latent-state trajectories. For instance, to sample a new state $\boldsymbol{x}_{t}$ given a new $\boldsymbol{\theta}$, one needs the full path
$p(\boldsymbol{x}_{0:t}\!$ $\mid\!\boldsymbol{\theta},h_{t-1})$
$\propto p(\boldsymbol{x}_{0})\prod_{s=1}^{t}f(\boldsymbol{x}_{s}\!\mid\!\boldsymbol{x}_{s-1},\boldsymbol{\theta},\boldsymbol{\xi}_{s})\,\prod_{s=1}^{t-1}g(\boldsymbol{y}_{s}\!\mid\!\boldsymbol{x}_{s},\boldsymbol{\theta},\boldsymbol{\xi}_{s}),$
which cannot be updated incrementally without revisiting all $s=1,\dots,t-1$. Naïvely, this yields $\mathcal{O}(t^{2})$ cost across $t$ steps, as each new gradient evaluation requires resimulating and reweighting entire trajectories.
In the next subsection, we show how this quadratic bottleneck can be avoided using the NPF framework, yielding Monte Carlo gradient estimators that require only a single forward propagation step per iteration, i.e., with complexity $\mathcal{O}(t)$.
3.3 Monte Carlo estimators using NPF
We now instantiate the terms in (3.2) using the approximation of the posterior distributions available at time t1t{-}1, obtained with the NPF framework. Let $\{\boldsymbol{\theta}_{t-1}^{(m)},\,w_{\boldsymbol{\theta},t-1}^{(m)}\}_{m=1}^{M}$ denote the outer (parameter) particles and, for each $m$, let $\{\boldsymbol{x}_{0:t-1}^{(m,n)},\,w_{\boldsymbol{x},t-1}^{(m,n)}\}_{n=1}^{N}$ be the inner (state-trajectory) particles. This yields particle approximations to $p(\boldsymbol{\theta}\!\mid\!h_{t-1})$ and $p(\boldsymbol{x}_{0:t-1}\!\mid\!\boldsymbol{\theta},h_{t-1})$ that we update online without reprocessing past data.
For approximating the EIG or any other quantity in (4) or (3.2) given a design $\boldsymbol{\xi}_{t}$, we first sample from the joint $\Gamma_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t},\boldsymbol{x}_{0:t},\boldsymbol{\theta})$. In the NPF framework, we reuse samples from the previous posterior and only draw one-step predictive states and pseudo-observations
$\widetilde{\boldsymbol{x}}_{t}^{(m,n)}$ $\sim f(\boldsymbol{x}_{t}\mid\boldsymbol{x}_{t-1}^{(m,n)},\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{\xi}_{t}),$
$\widetilde{\boldsymbol{y}}_{t}^{(m,n)}$ $\sim g(\boldsymbol{y}_{t}\mid\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{\xi}_{t}).$
Thus, Monte Carlo samples from $\Gamma_{\boldsymbol{\xi}_{t}}$ form the array $\{(\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{x}_{0:t-1}^{(m,n)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\widetilde{\boldsymbol{y}}_{t}^{(m,n)})\}$, needed for the outer expectation in (3.2), with weights
$w_{\boldsymbol{y},t}^{(m,n)}=w_{\boldsymbol{\theta},t-1}^{(m)}\,w_{\boldsymbol{x},t-1}^{(m,n)}.$
For notational clarity, we relabel the pseudo-observations as $\widetilde{\boldsymbol{y}}_{t}^{(\ell)}$, $\ell=1,\dots,L$, $L=M\times N$, with corresponding weights $w_{\boldsymbol{y},t}^{(\ell)}$, where each index $\ell$ corresponds to a unique pair $(m,n)$.
The samples required for the inner expectations that integrate out latent states are sampled in a similar way. We then sample the next state in the trajectory with different conditioning: (i) for $L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\widetilde{\boldsymbol{y}}_{t})$ we fix the parameters, $\boldsymbol{\theta}_{t-1}^{(m)}$; and (ii) for $Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ we sample new parameters in order to draw new states. With the NPF particles, we can operate:
(a) For $L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$: fix $\boldsymbol{\theta}=\boldsymbol{\theta}_{t-1}^{m}$ and propagate the associated inner particles one step,
$\ddot{\boldsymbol{x}}_{t}^{(m,j)}\sim f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1}^{(m,j)},\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{\xi}_{t}),$
to obtain the Monte Carlo approximation
$\widehat{L}_{\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{\xi}_{t}}^{N}(\boldsymbol{y}_{t})=\sum_{j=1}^{N}w_{\boldsymbol{x},t-1}^{(m,j)}\,g(\boldsymbol{y}_{t}\!\mid\!\ddot{\boldsymbol{x}}_{t}^{(m,j)},\boldsymbol{\theta}_{t-1}^{(m)},\boldsymbol{\xi}_{t}).$ (6)
(b) For $Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$: jitter parameter particles and propagate one step,
$\dot{\boldsymbol{\theta}}_{t-1}^{(i)}$ $\sim\kappa_{M}(\cdot\!\mid\!\boldsymbol{\theta}_{t-1}^{(i)}),$
$\dot{\boldsymbol{x}}_{t}^{(i,j)}$ $\sim f(\boldsymbol{x}_{t}\!\mid\!\boldsymbol{x}_{t-1}^{(i,j)},\dot{\boldsymbol{\theta}}_{t-1}^{(i)},\boldsymbol{\xi}_{t}),$
then average across both indices with the outer and inner weights,
$\widehat{Z}_{\boldsymbol{\xi}_{t}}^{M,N}(\boldsymbol{y}_{t})=\sum_{i=1}^{M}\sum_{j=1}^{N}w_{\boldsymbol{\theta},t-1}^{(i)}w_{\boldsymbol{x},t-1}^{(i,j)}\,g(\boldsymbol{y}_{t}\!\mid\!\dot{\boldsymbol{x}}_{t}^{(i,j)},\dot{\boldsymbol{\theta}}_{t-1}^{(i)},\boldsymbol{\xi}_{t}).$ (7)
Thus, to approximate the EIG in (4) we draw: (i) parameters $\{\boldsymbol{\theta}_{t}^{(m)},\dot{\boldsymbol{\theta}}_{t}^{(m)}\}$ for $m=1,\dots,M$; (ii) conditional state trajectories $\{\boldsymbol{x}_{0:t-1}^{(m,n)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\dot{\boldsymbol{x}}_{t}^{(m,n)},\ddot{\boldsymbol{x}}_{t}^{(m,n)}\}$ for $n=1,\dots,N$; and (iii) pseudo-observations $\widetilde{\boldsymbol{y}}_{t}^{(\ell)}\!\sim\!g(\boldsymbol{y}_{t}\!\mid\!\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}^{(m)},\boldsymbol{\xi}_{t})$, for $\ell=1,\dots,L$. This yields the nested Monte Carlo approximation
$\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})=\sum_{\ell=1}^{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{\boldsymbol{y},t-1}^{(\ell)}}\log\frac{\widehat{L}_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}^{N}(\widetilde{\boldsymbol{y}}_{t}^{(\ell)})}{\widehat{Z}_{\boldsymbol{\xi}_{t}}^{M,N}(\widetilde{\boldsymbol{y}}_{t}^{(\ell)})}.$ (8)
Following a similar procedure, the gradients $\nabla_{\boldsymbol{\xi}_{t}}L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ and $\nabla_{\boldsymbol{\xi}_{t}}Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ of (3.2) can be approximated. See details in Appendix LABEL:supp-ap:approximations.
3.4 Consistency of the EIG estimator
Consistency of our estimator $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ follows from two ingredients: (i) NPF convergence under standard conditions (small/regular jittering kernels, Lipschitz state posterior in $\boldsymbol{\theta}$, and a bounded positive likelihood), and (ii) NMC consistency provided the EIG integrand, $\log(L/Z)$, is Lipschitz and square-integrable. Formal statements of these assumptions are given in Appendix LABEL:supp-ap:consistency.
Theorem 1.
Let $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ denote the nested Monte Carlo estimator of the EIG defined in (8), constructed using $M$ parameter particles, $N$ state particles per parameter, and $L$ pseudo-observations. Under Assumptions 1–5 (in Appendix LABEL:supp-ap:consistency), it holds that
$\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})\;\xrightarrow[L,M,N\to\infty]{\text{a.s.}}\;\mathcal{I}(\boldsymbol{\xi}_{t})$
for any $t$ and $\boldsymbol{\xi}_{t}$.
The proof combines NMC results (Rainforth18), ensuring consistency of nested expectations under Assumption 5, with NPF convergence guarantees (Crisan18bernoulli), which establish almost-sure convergence of empirical state–parameter measures under Assumptions 1–4. Together, these results imply that the Monte Carlo averages in our estimator converge almost surely to the true EIG. Full details are given in Appendix LABEL:supp-ap:consistency.
3.5 Overall algorithm
Algorithm 1 summarizes our proposed framework, BAD-PODS (Bayesian Adaptive Design for Partially Observable Dynamical Systems). At each time step $t$, the design $\boldsymbol{\xi}_{t}$ is optimized by SGA (Section 3.2), using the gradient estimator of the EIG computed with NMC samples from $\Gamma_{\boldsymbol{\xi}_{t}}$ (Section 3.3). Iterative updates over $K$ steps, with step sizes $\eta_{k}$, yield the optimized design $\boldsymbol{\xi}_{t}^{\star}$, after which the real observation $\boldsymbol{y}_{t}$ is collected. The posterior is then updated via a NPF (Section 2.4): parameter particles are jittered, states are propagated through $f$, and weights are adjusted using $g$, with resampling to maintain diversity. This recursive construction avoids reprocessing past data, with a per-step cost of $\mathcal{O}((KL+1)MN)$ (and $\mathcal{O}(KLMNt)$ over $t$ steps), ensuring linear scaling in time while supporting gradient-based design optimization.
The algorithm takes as input the prior distributions, particle numbers $(M,N)$, and optimization hyperparameters $(K,\eta_{k})$, and produces as output the optimized design sequence $\{{\boldsymbol{\xi}}_{t}^{\star}\}_{t=1:T}$ together with posterior approximations of states and parameters. BAD-PODS therefore provides a practical realization of our framework, jointly optimizing designs and performing inference online, with consistency guarantees (Section 3.4) and support for continuous design spaces. Full implementation details are given in Appendix LABEL:supp-ap:algorithm.
Algorithm 1 BAD-PODS
1:Inputs: Number of particles $(L,M,N)$, stepsizes $\{\eta_{k}\}_{k=0}^{K-1}$, prior distributions $p(\boldsymbol{\theta})$, $p(\boldsymbol{x}_{0})$.
2:Init: $\boldsymbol{\theta}_{0}^{(m)}\!\sim\!p(\boldsymbol{\theta})$, $\boldsymbol{x}_{0}^{(m,n)}\!\sim\!p(\boldsymbol{x}_{0})$; set weights uniform.
3:for $t=1,2,\dots$ do
4:  /* Design optimization (SGA) */
5:  Initialize $\boldsymbol{\xi}_{t}^{(0)}$ randomly.
6:  for $k=0$ to K1K-1 do
7:   Sample $(\widetilde{\boldsymbol{y}}_{t}^{(\ell)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}_{t-1}^{(m)})$ from $\Gamma_{\boldsymbol{\xi}_{t}^{(k)}}$.
8:   Estimate $\widehat{L}_{\boldsymbol{\theta}_{t-1},\boldsymbol{\xi}_{t}^{(k)}}^{N}$ and $\widehat{Z}_{\boldsymbol{\xi}_{t}^{(k)}}^{M,N}$.
9:   Estimate EIG gradient of (3.2).
10:   Update $\boldsymbol{\xi}_{t}^{(k+1)}\leftarrow\boldsymbol{\xi}_{t}^{(k)}+\eta_{k}\,\widehat{\nabla_{\boldsymbol{\xi}_{t}}}\mathcal{I}(\boldsymbol{\xi}_{t}^{(k)})$.
11:  end for
12:  Set $\boldsymbol{\xi}_{t}^{\star}\leftarrow\boldsymbol{\xi}_{t}^{(K)}$; collect real $\boldsymbol{y}_{t}$.
13:
14:  /* Inference: update posterior distributions*/
15:  Jitter $\boldsymbol{\theta}_{t\mid t-1}^{(m)}\sim\kappa_{M}(\cdot\mid\boldsymbol{\theta}_{t-1}^{(m)})$.
16:  Propagate $\boldsymbol{x}_{t\mid t-1}^{(m,n)}\sim f(\cdot\!\mid\!\boldsymbol{x}_{t-1}^{(m,n)},\boldsymbol{\theta}_{t\mid t-1}^{(m)},\boldsymbol{\xi}_{t}^{\star})$.
17:  Compute weights with $g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t\mid t-1}^{(m,n)},\boldsymbol{\theta}_{t\mid t-1}^{(m)},\boldsymbol{\xi}_{t}^{\star})$.
18:  Resample first state particles, then parameter particles.
19:end for
4 RELATED WORK
A central challenge in BED is the cost and complexity of estimating the EIG (or its gradient) (Rainforth24; huan2024optimal). For models with tractable likelihoods $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, the nested Monte Carlo (NMC) estimator has been studied in depth (Rainforth18), with subsequent work proposing variational formulations and bounds to improve convergence (Foster19; Foster20; Foster21). These approaches also enable differentiable objectives for gradient-based design in continuous spaces. However, they rely on explicit pointwise likelihood evaluation and do not extend directly to partially observable dynamical systems, where likelihoods require marginalization over latent states.
When likelihoods are implicit (intractable pointwise but possible to simulate from), likelihood-free design methods replace direct evaluation by surrogates or simulation-based objectives. Examples include variational mutual-information bounds and amortized estimators (Kleinegesse20; Kleinegesse21; Ivanova21), density-ratio or classifier-based estimates of information gain (Kleinegesse19), and approximate Bayesian computation (ABC)-style utilities (Dehideniya18; Drovandi13; Hainy16; Price16). While effective in simulator settings, most such methods target static design (optimize once, then deploy) rather than fully sequential/adaptive BED. Moreover, surrogate-based adaptive variants (e.g., Ivanova21) typically require an offline dataset to train an implicit likelihood, which may be unavailable or quickly misspecified in evolving dynamical systems.
Design for dynamical systems has also been studied under full state observability. Iqbal24; iqbal2024recursive consider adaptive design when the state is fully observed, making the likelihood tractable. Both papers leverage particle methods for joint state–parameter inference: (Iqbal24) uses a reversed (“inside-out”) SMC2, and (iqbal2024recursive) adopts NPF’s jittering to rejuvenate parameter particles. However, these recent advances in BED for dynamical systems assume full state observability and do not address the latent-state marginalizations required in partially observable SSMs.
5 EXPERIMENTS
We evaluate our online BAD framework on two partially observable dynamical systems. Since no existing methods target this setting, we compare against two standard BED baselines (foster2021thesis): (i) random designs sampled uniformly from $\Omega$, and (ii) static BED, a non-adaptive version of our method where the full sequence $\boldsymbol{\xi}_{1:T}$ is optimized offline. Unlike our sequential approach, the static baseline cannot adapt as data arrive.
Performance is measured by the total EIG (TEIG), the accumulated information gain across time. We also report relative improvements,
$\Delta\text{TEIG}^{\text{(baseline)}}=\sum_{s=1}^{t}\bigg(\widehat{\mathcal{I}}(\boldsymbol{\xi}_{s}^{\text{BAD-PODS}})-\widehat{\mathcal{I}}(\boldsymbol{\xi}_{s}^{\text{baseline}})\bigg),$
against each baseline to highlight adaptive gains. Results are averaged over 50 seeds with bootstrap 95% confidence intervals.
Python code to reproduce all experiments is released at https://anonymous.4open.science/r/badpods-2C8D. Details of the computing infrastructure are provided in Appendix LABEL:supp-ap:infrastructure.
5.1 Two-group SIR model
The sequential importance resampling (SIR) model partitions a population into susceptible ($S$), infected ($I$), and recovered ($R$) compartments. It has been widely used as a testbed in BED, though prior work often assumes fully observed states (Ivanova21) or static, likelihood-free design (Kleinegesse19).
We instead consider a partially observable, stochastic two-group SIR with online sequential design. Each group $g\in\{1,2\}$ has latent states $\mathbf{X}^{g}(\tau)=(S^{g}(\tau),I^{g}(\tau))^{\top}$ (the recovered $R^{g}$ being redundant since total population is fixed to $N_{g}=S^{g}+I^{g}+R^{g}$). The model parameters are infection and recovery rates $\boldsymbol{\theta}=\{(\beta_{g},\gamma_{g})\}_{g=1}^{2}$, with cross-group mixing governed by a fixed matrix $M\in\mathbb{R}^{2\times 2}$.
Dynamics follow an Itô stochastic differential equation (SDE) (allen2008introduction; ArinoVanDenDriessche2003; chatzilena2019contemporary) of the form,
$\mathrm{d}\mathbf{X}=f(\mathbf{X})\,\mathrm{d}\tau+G(\mathbf{X})\,\mathrm{d}\mathbf{W},$
with drift $f$, diffusion matrix $G$, and independent Wiener processes $\mathbf{W}$.
At each discrete time $t$, a design $\boldsymbol{\xi}_{t}=(\xi_{t}^{(1)},\xi_{t}^{(2)})$ allocates sampling effort ($\kappa>0$) across groups, constrained to the 2-simplex. Observations are noisy incident counts,
$y_{t}^{(g)}\;\big|\;\mathbf{X}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t}\;\sim\;\mathrm{Poisson}\!\left(\kappa\,\xi_{t}^{(g)}\,\rho_{g}\,\frac{I_{t}^{g}}{N_{g}}\right),$ (9)
with known detection scales $\rho_{g}>0$. Thus $\boldsymbol{\xi}_{t}$ affects only the observation model. Simulation is performed using an Euler–Maruyama discretization. Full model details are given in Appendix LABEL:supp-ap:sir.
Experimental setup.
We set the horizon to $T=200$ and simulate ground-truth trajectories from the two-group SIR model with group sizes $N_{1}\!=\!N_{2}\!=\!200$ and initial infections $I_{0}^{1}\!=\!I_{0}^{2}\!=\!5$. The cross-group mixing matrix $M$, detection scales $\rho_{g}$, sampling budget $\kappa$, and the parameters of the second group $(\beta_{2},\gamma_{2})$ are fixed and known. Inference targets only the first-group parameters $(\beta_{1},\gamma_{1})$. At each time step, the design $\boldsymbol{\xi}_{t}$ allocates observation effort between $y_{t}^{(1)}$ and $y_{t}^{(2)}$.
Our online BAD framework is implemented with a NPF using M=N=100M\!=\!N\!=\!100 particles. Design optimization at each step uses K=500K\!=\!500 stochastic gradient updates, with simplex constraints enforced via a sigmoid reparameterization. Updates are performed with Adam (DBLP:journals/corr/KingmaB14). Further details are given in Appendix LABEL:supp-ap:sir.
Experimental results.
Table 1 reports TEIG for all three methods. BAD-PODS consistently outperforms both baselines, with gaps widening over time. Results for the static method are truncated as optimization becomes infeasible at longer horizons.
Table 1: TEIG across time for two-group SIR (left) and moving source (right). Means over 50 seeds with 95% BCa CIs. Static results truncated at longer horizons due to computational cost.
Two-group SIR model Moving source model
$t$ BAD-PODS Random Static $t$ BAD-PODS Random Static
50 1.390 [1.324,1.460] 1.364 [1.287,1.445] 1.167 [1.117,1.220] 10 0.304 [0.290,0.325] 0.242 [0.222,0.263] 0.270 [0.258,0.284]
100 3.283 [3.173,3.380] 3.152 [3.040,3.268] 2.529 [2.444,2.621] 20 0.709 [0.650,0.855] 0.582 [0.531,0.649] 0.570 [0.541,0.604]
150 4.748 [4.609,4.871] 4.518 [4.385,4.627] 3.547 [3.454,3.658] 30 1.136 [1.005,1.370] 0.940 [0.855,1.044]
200 5.918 [5.765,6.055] 5.621 [5.466,5.793] 40 1.483 [1.318,1.822] 1.217 [1.136,1.344]
50 1.740 [1.551,2.096] 1.435 [1.354,1.555]
[…]
Figure 1: Two-group SIR: differences in TEIG (ours minus baseline). Boxplots (median and interquartiles) compare BAD-PODS with random designs (left, green) and static BED (right, orange). Higher values indicate better performance for BAD-PODS.
Figure 2: Distribution of design component $\xi_{t}^{(1)}$ across seeds and time (boxplots: median and interquartiles). BAD-PODS allocates more effort to group 1, which contains the unknown parameters.
5.2 Moving source location
Source localization is a standard testbed in BED (Ivanova21; Foster21), usually formulated with a static source and sensor positions as designs. Here we consider a more challenging variant with a moving source, where the designs are the orientations of fixed sensors.
We model a single source with latent state $\boldsymbol{x}_{t}\!=\!(p_{x,t},p_{y,t},\phi_{t})^{\top}\in\mathbb{R}^{2}\!\times\!(-\pi,\pi]$ and motion parameters $\boldsymbol{\theta}\!=\!(v_{x},v_{y},v_{\phi})$ (full dynamics in Appendix LABEL:supp-ap:source). $J$ sensors are located at fixed positions $\{\boldsymbol{s}_{j}\}_{j=1}^{J}\subset\mathbb{R}^{2}$, and at each time step $t$ the sensor orientations $\boldsymbol{\xi}_{t}\!=\!(\xi_{t,1},\ldots,\xi_{t,J})\in[-\pi,\pi)^{J}$ need to be specified. Each sensor has a directional response: the effective intensity depends on the angular mismatch $\Delta_{t,j}\!=\!\xi_{t,j}-\psi_{t,j}(\boldsymbol{p}_{t})$, where $\psi_{t,j}(\boldsymbol{p}_{t})=\operatorname{atan2}\!\big((\boldsymbol{p}_{t}-\boldsymbol{s}_{j})_{y},(\boldsymbol{p}_{t}-\boldsymbol{s}_{j})_{x}\big)$ is the bearing to the source.
The observation model follows (sheng2005maximum). Each sensor reports a noisy log–intensity,
$\log y_{t,j}$ $\mid\boldsymbol{p}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t}\sim\mathcal{N}\!\big(\log\mu_{t,j},\,\sigma^{2}\big),$ (10)
$\mu_{t,j}$ $=b\;+\;\frac{\alpha_{j}}{\,m+\|\boldsymbol{p}_{t}-\boldsymbol{s}_{j}\|^{2}\,}\,D\!\big(\Delta_{t,j}\big),$ (11)
with cardioid directivity $D(\delta)=\big(\frac{1+d\cos\delta}{1+d}\big)^{k}$ ($d\in[0,1)$), background $b>0$, saturation $m>0$, directional strength exponent $k>0$ and sensor strengths $\alpha_{j}\geq 0$. The design $\boldsymbol{\xi}_{t}$ affects only the observation model.
Experimental setup.
We simulate a moving source over a horizon of T=50T\!=\!50 steps starting at position $\boldsymbol{x}_{0}\!=\!(0,0,0)$. The source parameters to be inferred are the horizontal and vertical velocities, $\boldsymbol{\theta}=(v_{x},v_{y})$, while the angular velocity $v_{\phi}$ and all other observation parameters ($b$, $m$, $d$, $k$, $\alpha_{j}$, $\sigma^{2}$) are treated as known. We deploy J=2J\!=\!2 fixed sensors located at $\boldsymbol{s}_{1}\!=\!(0,3)$ and $\boldsymbol{s}_{2}\!=\!(3,0)$, each with orientation design variables $\boldsymbol{\xi}_{t}\!=\!(\xi_{t,1},\xi_{t,2})\in[-\pi,\pi)^{2}$. Figure 3 illustrates the setup and one example trajectory. Inference is performed with a NPF using M=N=300M\!=\!N\!=\!300 particles, and design optimization at each step uses K=300K\!=\!300 stochastic gradient updates. Optimization is carried out with Adam (DBLP:journals/corr/KingmaB14). Additional implementation details and parameter values are given in Appendix LABEL:supp-ap:source.
Figure 3: Example trajectory of the moving source for t=50t\!=\!50 (blue line with markers) together with fixed sensor locations (red triangles).
Experimental results.
Table 1 reports the total EIG (TEIG) for our sequential method compared against the random and static baselines. Once again, our approach achieves the highest TEIG across all horizons, confirming the benefits of adapting designs online in partially observable settings. Figure 4 illustrates these results more clearly by showing boxplots of TEIG differences relative to each baseline. The improvements accumulate steadily over time as information gain compounds. As in the SIR experiment, results for the static method are truncated at larger horizons due to computational infeasibility of offline optimization in high-dimensional design spaces.
Figure 4: Moving source location: differences in TEIG (ours minus baseline). Boxplots (median and interquartiles) compare BAD-PODS with random designs (left, green) and static BED (right, orange). Higher values indicate better performance for BAD-PODS.
Finally, we assess the quality of the selected sensor orientations in terms of their pointing error, i.e., the angular deviation between a sensor’s chosen orientation and the true bearing to the source. Figure 5 shows boxplots of pointing error distributions at selected time steps across seeds and sensors. Random designs exhibit wide variability, often pointing far away from the target. The static baseline is more stable, but it still accumulates errors over time, as it cannot adapt to the evolving state of the system. In contrast, our sequential method consistently achieves much smaller pointing errors, demonstrating its ability to align sensors toward the target more frequently and with greater accuracy.
Figure 5: Pointing error (in degrees) of sensor orientations relative to the target at selected time steps. Boxplots (median and interquartiles) show the distribution across seeds and sensors for our method, random design, and static BED approach. Lower values correspond to higher orientation accuracy.
6 CONCLUSIONS
We introduced an online Bayesian adaptive design (BAD) framework for partially observable dynamical systems, modeled as state-space models (SSMs). Our approach combines two key ingredients: (i) new estimators of the expected information gain and its gradient that explicitly account for latent states, and (ii) an nested particle filter (NPF)-based construction that reuses state–parameter particles to evaluate these quantities online. This yields an algorithm whose cost grows only linearly with time, avoids reprocessing past data, and remains asymptotically consistent under mild assumptions. The framework supports continuous design spaces and stochastic-gradient optimization, making it broadly applicable. Through experiments on a multi-group epidemiological model and a dynamical source–tracking problem, we demonstrated that online adaptation significantly improves information efficiency over static Bayesian experimental design (BED) or random design baselines.